Suppose venture capital form GSB partners raised 100 million of committed capital. Each year over the 9 year life of the fund, 2.5% of this committed capital will be used to pay GSB's management fee. As is typical in the venture capital industry. GSB will only invest 77.50 million (committed capital less lifetime management fees). At the end of 9 years, the investment made by the fund is worth 600 million. GSB also charges 30% carried interest on the profit of the fund (net of management fees).

A. Assume the 77.50 million of invested capital is invested immediately, and all proceeds were received at the end of 9 years. What is the IRR of the investments GSB partners made?

B. Of course, as an investor or limited partner, you are more interested in your own IRR, that is the IRR including all fees paid. Assuming that investors gave GSB partners made? What is the IRR for GSB's limited partners (that is, the IRR net of all fees paid)?
Answer #1

A.

Invested Amount at time t=0 = 77.5 million (cash outflow)

Investment Value at end of nine year, t=9 = 600 million (cash inflow)

Therefore, IRR = r%, then 77.5 = 600 / (1+r)^9

Solving for r, r= 25.53%.

Therefore IRR of investment made is 25.53%

B.

Invested Amount at time t=0 = 77.5 million (Cash Outflow)

For t=1 to t=8 , cash outflow = 2.5 million each year

For t=9, cash outflow towards management fees = 2.5 million

investment amount = 600 million

profit = 600-2.5 = 597.5

carried interest = 0.3*597.5 = 179.25

Therefore, cash inflow for investor = 600-2.5-179.25 = 418.25

Solving for IRR, IRR = 19.00%